### v1

# On dense subalgebras of the singular ideal in groupoid C*-algebras

Abstract.
We prove that ideals in amenable second-countable non-Hausdorff étale groupoid $C^{*}$-algebras are determined by their isotropy fibres. As an application, we characterise when the singular functions in Connes’ algebra are dense in the singular ideal in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras.
Key words and phrases:
Étale groupoid, non-Hausdorff, $C^{*}$-algebra, singular ideal, isotropy fibres
2020 Mathematics Subject Classification:
22A22, 46L05 (Primary); 37A55, 22D25 (Secondary)
Introduction
In recent years non-Hausdorff groupoids and their $C^{*}$-algebras have gained increased attention. Many groupoids arising from dynamics and geometry are non-Hausdorff, and it is therefore important to develop a robust theory. Indeed, there are natural examples of groupoids of germs, groupoids arising from self-similar groups, and groupoids arising from foliations which are non-Hausdorff ([10], [21], [22], [23]).
In contrast with the Hausdorff case, the functions in the reduced $C^{*}$-algebra $C^{*}_{r}(G)$ of a non-Hausdorff groupoid $G$ are not necessarily continuous. Indeed, there can exist (non-zero) functions in $C^{*}_{r}(G)$ whose set of non-zero values has empty interior in $G$. The set of all such functions is known as the singular ideal $J$ in $C^{*}_{r}(G)$. Historically, the singular ideal has been an obstruction to understanding simplicity for reduced $C^{*}$-algebras of non-Hausdorff groupoids ([9]). Characterisations of simplicity and the ideal intersection property have been obtained for the quotient by this ideal ([9], [18]), which is known as the essential groupoid $C^{*}$-algebra ([19], [13]). Therefore, the only obstacle to understanding these properties for the reduced $C^{*}$-algebra is the singular ideal.
Much work has been done to find conditions that ensure the singular ideal $J$ vanishes (see [9], [19], [24], [15]). Recently in [16] vanishing of the singular ideal was characterised in terms of a property of certain collections of subgroups in isotropy groups of $G$. In [4] it is characterised when the singular ideal $J$ has trivial intersection with the underlying groupoid -algebra, $\mathscr{C}_{c}(G)$ known as Connes’ algebra (first defined in [10] for groupoids arising from foliations).
When the singular ideal is non-zero, less is known about its structure. For example, it is not known in general whether $J\neq\{0\}$ implies $J\cap\mathscr{C}_{c}(G)\neq\{0\}$. This was shown for groupoids with finite “non-Hausdorfness” in [4] and in [16] the problem was shown to be equivalent to whether group $C^{*}$-algebras satisfy the Intersection Property. This reduction of the problem led to its solution for groupoids with isotropy groups that are direct limits of virtually torsion free solvable groups [16].
A related question asks for which groupoids is $J\cap\mathscr{C}_{c}(G)$ dense in $J$ ([4, Question 4.11(III)]). Progress could be made towards understanding the structure of $J$ if it has a dense subalgebra of functions in $\mathscr{C}_{c}(G)$ (since functions in $\mathscr{C}_{c}(G)$ are easier to understand). In recent work ([20]), it is shown that there exist non-amenable groupoids for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$. The following question remains open however.
Question.
Is $J\cap\mathscr{C}_{c}(G)$ dense in $J$ for any amenable étale groupoid $G$?
The goal of our paper is to study the question highlighted above. Let us outline the main achievements of the paper. All results listed here are for amenable second-countable étale groupoids.
  • We show that ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres. This generalises [8, Theorem 2.10] to the setting of non-Hausdorff groupoids.
  • Following the approach in [16], we characterise when $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras.
  • We describe some large classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$. In particular, we show density holds for groupoids with abelian isotropy groups, and groupoids arising from contracting self-similar groups.
We now state our main results and describe them in more detail. Our first result concerns ideals in groupoid $C^{*}$-algebras. In [8] it is shown that ideals $I\trianglelefteq C^{*}(G)$ in a groupoid $C^{*}$-algebra are mapped to ideals in the isotropy group $C^{*}$-algebras by the restriction maps $C^{*}(G)\to C^{*}(G^{x}_{x})$. These images are known as the isotropy fibres of the ideal, and are denoted $(I_{x})_{x\in G^{(0)}}$. It is shown in [8, Theorem 2.10] that for amenable second-countable Hausdorff étale groupoids, ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres. We generalise this result to non-Hausdorff groupoids.
Theorem A (See Theorem 2.5).
Let $G$ be an amenable and second-countable (non-Hausdorff) étale groupoid and suppose $I$ and $K$ are ideals in $C^{*}(G)$. Then, $I=K$ if and only if $I_{x}=K_{x}$ for all $x\in G^{(0)}$.
We apply this result to the singular ideal $J$ in the reduced $C^{*}$-algebra of a non-Hausdorff groupoid. The following Theorem reduces the density question for $J\cap\mathscr{C}_{c}(G)$ in $J$ to one about density of ideals in isotropy group $C^{*}$-algebras.
Theorem B (See Proposition 2.6).
Let $G$ be an amenable and second-countable étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ if and only if $J_{x}\cap\mathbb{C}[G^{x}_{x}]$ is dense in $J_{x}$ for all $x\in G^{(0)}$.
The isotropy fibres $J_{x}$ are calculated explicitly in [16, Theorem 5.5]. Moreover, when $G$ is covered by countably many bisections and $G^{x}_{x}$ is amenable, it is shown $J_{x}$ equals the ideal $J_{G^{x}_{x},\mathcal{X}(x)}$ defined as the intersection of the kernels of quasi-regular representations associated to subgroups of $G^{x}_{x}$ in a certain a collection $\mathcal{X}(x)$. We refer the reader to Subsection 2.1 for the definition of $\mathcal{X}(x)$ and $J_{G^{x}_{x},\mathcal{X}(x)}$. The following is then a rephrasing of Theorem B.
Theorem C (See Theorem 2.9).
Let $G$ be an amenable and second-countable étale groupoid. Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ if and only if $J_{G^{x}_{x},\mathcal{X}(x)}\cap\mathbb{C}[G^{x}_{x}]$ is dense in $J_{G^{x}_{x},\mathcal{X}(x)}$ for all $x\in G^{(0)}$. In particular, if all the isotropy groups $G^{x}_{x}$ satisfy the Density Property, then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
In the above theorem, the Density Property is a property of discrete amenable groups we introduce in Definition 2.8. Finite groups satisfy the Density Property automatically, and in Proposition 3.11 we show that all abelian groups have the Density Property. Currently, we do not know of any discrete amenable group that fails the Density Property. We remark that it is a corollary of Theorem C and a non-Hausdorff groupoid construction in [16, Section 6] that if there is such a group, then there is an amenable groupoid for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$ (see Corollary 2.12).
By the definition of the ideals $J_{G^{x}_{x},\mathcal{X}(x)}$, Theorem C asserts that density of $J\cap\mathscr{C}_{c}(G)$ in $J$ can be understood entirely in terms of quasi-regular representations of isotropy groups. In light of this, we are able to describe some explicit classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Theorem D (See Corollary 3.12).
Let $G$ be an amenable and second-countable étale groupoid. Assume for every $x\in G^{(0)}$ one of the following holds.
  1. (I)
    The isotropy group $G_{x}^{x}$ is abelian.
  2. (II)
    The subgroups $X\in\mathcal{X}(x)$ are finite.
Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Theorem D covers many important classes of groupoids. In particular, any groupoid of germs associated with the action of a countable abelian group will satisfy (I). All groupoids with finite isotropy groups will satisfy (II). More generally, condition (II) is satisfied whenever the unit space of $G$ has finite source and range fibres.
In the setting of non-Hausdorff ample groupoids, one can intersect $J$ with the complex Steinberg algebra $\mathbb{C}G$ (see [26]) to obtain what is known as the algebraic singular ideal $J_{\mathbb{C}}$. A characterisation for the vanishing of $J_{\mathbb{C}}$ is given in [4], and together with [9] and [27] this leads to a complete characterisation of simplicity for Steinberg algebras (see also [28] and [15]). In this setting, one can ask the following density question: for which ample groupoids is $J_{\mathbb{C}}$ dense in $J$? Indeed, this is presented as an open problem in [15]. In order to answer this question, we show it suffices to study density of $J\cap\mathscr{C}_{c}(G)$ in $J$.
Theorem E (See Theorem 4.4).
Let $G$ be an ample groupoid. Then $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
An important class of ample groupoids are those arising from self-similar groups (see [22]). These groupoids can be non-Hausdorff, and can also have non-vanishing singular ideal. Using our results, we answer the density question for groupoids arising from contracting self-similar groups.
Corollary F (See Corollary 4.5).
Let $G$ be the groupoid arising from the action of a contracting self-similar group on a finite alphabet. Then $J_{\mathbb{C}}$ is dense in $J$.
In [20], examples of groupoids arising from self-similar actions on infinite alphabets are described for which $J_{\mathbb{C}}$ is not dense in $J$.
Acknowledgments
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Topological groupoids and their $C^{*}$-algebras, where work on this paper was undertaken. This work was supported by EPSRC grant EP/V521929/1. Hume was supported by NSERC Discovery Grant RGPIN-2021-03834. Gonzales would like to thank Xin Li for helpful discussions on ideas related to this paper.
1. Preliminaries
1.1. Étale groupoids
[…]
A topological groupoid $G$ is a groupoid equipped with a topology such that the product and inverse maps are continuous. By an étale groupoid $G$, we mean a locally compact topological groupoid for which the unit space $G^{(0)}$ is Hausdorff, and the range $r\colon G\to G^{(0)}$ and source $s\colon G\to G^{(0)}$ maps are local homeomorphisms. This paper will only consider étale groupoids, and primarily concerns non-Hausdorff groupoids (all results also hold for Hausdorff groupoids, but often for trivial reasons i.e. because the singular ideal vanishes for Hausdorff groupoids).
An open subset $U\subseteq G$ will be called an open bisection if the restrictions $r|_{U}\colon U\to r(U)$ and $s|_{U}\colon U\to s(U)$ are homeomorphisms. Note that any étale groupoid has a basis of open bisections. We define
$\mathscr{C}_{c}(G)\coloneq\text{span}\{f\colon G\to\mathbb{C}\colon f|_{U}\in C_{c}(U)\text{ and }f|_{G\setminus U}=0\text{ for some open bisection }U\subseteq G\}.$
If the groupoid $G$ is Hausdorff, then $\mathscr{C}_{c}(G)=C_{c}(G)$, the set of all compactly supported continuous functions on $G$. However if $G$ is not Hausdorff, then functions in $\mathscr{C}_{c}(G)$ need not be continuous. We equip $\mathscr{C}_{c}(G)$ with the structure of a -algebra as follows. For $f,f_{1},f_{2}\in\mathscr{C}_{c}(G)$, and $g\in G$ define
$f^{*}(g)\coloneq\overline{f(g^{-1})}\quad\text{and}\quad f_{1}*f_{2}(g)\coloneq\sum_{h\in G_{s(g)}}f_{1}(gh^{-1})f_{2}(h).$
We now describe the reduced and full groupoid $C^{*}$-algebras. Given $x\in G^{(0)}$, let $\lambda_{x}\colon\mathscr{C}_{c}(G)\to B(\ell^{2}(G_{x}))$ be the representation given by $\lambda_{x}(f)(\xi)=f*\xi$ for $f\in\mathscr{C}_{c}(G)$ and $\xi\in\ell^{2}(G_{x})$, where $(f*\xi)(g)=\sum_{h\in G_{x}}f(gh^{-1})\xi(h)$ for $g\in G_{x}$. Then the reduced groupoid $C^{*}$-algebra $C^{*}_{r}(G)$ of $G$ is defined to be the completion of $\mathscr{C}_{c}(G)$ with respect to the norm $\|f\|_{C^{*}_{r}(G)}=\sup_{x\in G^{(0)}}\|\lambda_{x}(f)\|$. The full groupoid $C^{*}$-algebra $C^{*}(G)$ of $G$ is defined to be the completion of $\mathscr{C}_{c}(G)$ with respect to the norm $\|f\|_{C^{*}(G)}=\sup_{\pi}\|\pi(f)\|$, where the supremum is taken over all -algebra representations $\pi\colon\mathscr{C}_{c}(G)\to B(H)$, for some Hilbert space $H$.
Elements of the reduced groupoid $C^{*}$-alegbra $C^{*}_{r}(G)$ can be viewed as bounded Borel functions on $G$ in a canonical way (see [6, Lemma 3.27], [25, Proposition II.4.2]). From now on, we will always treat elements of $C^{*}_{r}(G)$ as functions on $G$ in this way. Given $a\in C^{*}_{r}(G)$, we define
${\rm supp}^{\circ}(a)\coloneq\{g\in G\colon a(g)\neq 0\}.$
Definition 1.1 ([9], [19], [13]).
The singular ideal $J$ in $C^{*}_{r}(G)$ is defined as
$J\coloneq\{a\in C^{*}_{r}(G)\colon{\rm supp}^{\circ}(a)\text{ has empty interior}\}.$
The singular ideal $J$ is a (closed) ideal in $C^{*}_{r}(G)$. For $a\in C^{*}_{r}(G)$, ${\rm supp}^{\circ}(a)$ has empty interior in $G$ if and only if $s\big({\rm supp}^{\circ}(a)\big)$ has empty interior in $G^{(0)}$ (see [1, Proposition 4.6], [4, Lemma 4.1(iii)]).
1.2. Hausdorff cover
We now introduce the Hausdorff cover groupoid. See [4] for an extensive study of the Hausdorff cover for non-Hausdorff étale groupoids. Let $G$ be a (non-Hausdorff) étale groupoid, and let $\mathfrak{C}(G)$ be its space of closed subsets. Singleton sets in $G$ are closed, hence there is a canonical inclusion $\iota\colon G\to\mathfrak{C}(G)$ given by $\iota(g)\coloneq\{g\}$. When equipped with the Fell topology, $\mathfrak{C}(G)$ has the structure of a compact Hausdorff space (see [14]).
Definition 1.2 ([29]).
Let $G$ be an étale groupoid. The Hausdorff cover $\tilde{G}$ is defined to be the closure of $\iota(G)$ in $\mathfrak{C}(G)\setminus\{\emptyset\}$ with respect to the Fell topology.
Elements of $\tilde{G}$ can be described explicitly as follows. A non-empty subset $\bm{g}\subseteq G$ lies in $\tilde{G}$ if and only if there exist a net $(g_{i})$ in $G$ such that $\bm{g}$ is the set of limit points of $(g_{i})$, and every limit point of a subnet of $(g_{i})$ (cluster point) is a limit point of $(g_{i})$. Moreover, given a net $(g_{i})$ in $G$, the image net $\iota(g_{i})$ converges to $\bm{g}\in\tilde{G}$ in the topology of $\tilde{G}$ if and only if the conditions above are satisfied.
Define $\tilde{G}^{(0)}\coloneq\{\bm{g}\in\tilde{G}\colon\bm{g}\cap G^{(0)}\neq\emptyset\}$. This set will be the unit space of $\tilde{G}$ with the structure described below. The canonical inclusion map $\iota\colon G\hookrightarrow\tilde{G}$ restricts to an inclusion $G^{(0)}\hookrightarrow\tilde{G}^{(0)}$, and the image $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$. For each $\bm{x}\in\tilde{G}^{(0)}$ there is a unique element $\pi(\bm{x})\in\bm{x}\cap G^{(0)}$, and the map $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is a continuous surjection. Since $\iota(G^{(0)})$ is dense in $\tilde{G}^{(0)}$, continuity of the groupoid operations on $G$ implies that any element $\bm{x}\in\tilde{G}^{(0)}$ is a subgroup of the isotropy group $G_{\pi(\bm{x})}^{\pi(\bm{x})}$. We will often write $\bm{x}=X$ to distinguish elements $x\in G^{(0)}$ from $\bm{x}\in\tilde{G}^{(0)}$.
We now describe the groupoid operations on $\tilde{G}$. Note that $r(g_{1})=r(g_{2})$ and $s(g_{1})=s(g_{2})$ whenever $g_{1},g_{2}\in\bm{g}$ and $\bm{g}\in\tilde{G}$. The fact that elements of $\tilde{G}^{(0)}$ are subgroups implies, for every $\bm{g}\in\tilde{G}$, there are unique elements $\bm{x},\bm{y}\in\tilde{G}^{(0)}$ such that $\bm{g}=g\cdot\bm{x}=\bm{y}\cdot g$ for any $g\in\bm{g}$. Let $s(\bm{g})=\bm{x}$ and $r(\bm{g})=\bm{y}$. Define the inverse $\bm{g}^{-1}\coloneq\{g^{-1}\colon g\in\bm{g}\}$ and the product $\bm{g}\bm{h}\coloneq\{gh\colon g\in\bm{g},h\in\bm{h}\}$ whenever $s(\bm{g})=r(\bm{h})$. Note also that
$\bm{g}\bm{h}=\{gh\colon h\in\bm{h}\}\text{ for any }g\in\bm{g}.$ (1.1)
With this structure, $\tilde{G}$ becomes a Hausdorff étale groupoid.
[…]
$\mathfrak{i}(f)(\bm{g})=\sum_{g\in\bm{g}}f(g)$ (1.2)
for all $\bm{g}\in\tilde{G}$ ([4, Lemma 3.2]). The embedding $Cc(G~)\mathfrak{i}\colon\mathscr$$\hookrightarrow C_{c}(\tilde{G})$ is a -homomorphism, and induces canonical embeddings $Cr(G~)\mathfrak{i}_{r}\coloneq C^{*}_{r}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}_{r}(\tilde{G})$ and $C(G~)\mathfrak{i}\coloneq C^{*}(\mathfrak{i})\colon$ $\hookrightarrow C^{*}(\tilde{G})$ ([4, Lemma 3.8 & Corollary 6.8]).
Now assume that the groupoid $G$ is covered by countably many open bisections. We say a unit $x\in G^{(0)}$ is Hausdorff if $\pi^{-1}(x)=\{\iota(x)\}$, and we let $C\subseteq G^{(0)}$ denote the set of Hausdorff units. By [19, Lemma 7.15], $C$ is dense in $G^{(0)}$ (the Hausdorff points complement the so-called dangerous points of [19]). Define $\tilde{G}^{(0)}_{\rm{ess}}\coloneq\overline{\iota(C)}$. This is an invariant subset of the unit space $\tilde{G}^{(0)}$. We define the essential Hausdorff cover $\tilde{G}_{\rm{ess}}\coloneq\tilde{G}|_{\tilde{G}^{(0)}_{\rm{ess}}}=\{\bm{g}\in\tilde{G}\colon s(\bm{g})\in\tilde{G}^{(0)}_{\rm{ess}}\}$ ([4, Definition 4.13]).
1.3. Isotropy fibres of ideals
Our attention now turns to restriction maps. Let $G$ be an étale groupoid. Given a unit $x\in G^{(0)}$, consider the restriction map $\eta_{x}\colon\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$, where $\mathbb{C}[G^{x}_{x}]$ denotes the group algebra of the isotropy group $G^{x}_{x}$ (thought of as finitely supported functions $G^{x}_{x}\to\mathbb{C}$). The map $\eta_{x}$ extends to a completely positive contraction $C^{*}(G)\to C^{*}(G^{x}_{x})$, which we will also denote by $\eta_{x}$. This can be seen using induced representations of the isotropy group (see [7, Lemma 1.2] for example). In fact, any $C^{*}$-norm on $\mathscr{C}_{c}(G)$ dominating the reduced norm will define a $C^{*}$-norm on the group algebra $\mathbb{C}[G^{x}_{x}]$ dominating the reduced norm, and the restriction map $\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$ will always extend to a completely positive contraction between the associated $C^{*}$-algebras. See [8, Section 1.3] for a discussion in this direction. For a (closed) ideal $I\trianglelefteq C^{*}(G)$, the image $\eta_{x}(I)$ is also a (closed) ideal in $C^{*}(G^{x}_{x})$ ([8, Lemma 2.1]). We write $I_{x}\coloneq\eta_{x}(I)$ and, following [8], call $I_{x}$ the isotropy fibre of $I$ at $x$.
For $g\in G^{x}_{x}$, denote by $\delta_{g}\in\mathbb{C}[G^{x}_{x}]$ the function equal to one at $g$ and zero everywhere else. Each $g\in G$ induces an isomorphism of group algebras $\text{Ad}_{g}\colon\mathbb{C}\big[G^{s(g)}_{s(g)}\big]\to\mathbb{C}\big[G^{r(g)}_{r(g)}\big]$, $\delta_{h}\mapsto\delta_{ghg^{-1}}$ and this maps extends to an isomorphism of $C^{*}$-algebras $\Psi_{g}\colon C^{*}\big(G^{s(g)}_{s(g)}\big)\to C^{*}\big(G^{r(g)}_{r(g)}\big)$. A family $(K_{x})_{x\in G^{(0)}}$ of ideals satisfying $K_{x}\trianglelefteq C^{*}(G_{x}^{x})$ for each $x\in G^{(0)}$ is said to be invariant if $\Psi_{g}(K_{s(g)})=K_{r(g)}$ for all $g\in G$. Given an ideal $I\trianglelefteq C^{*}(G)$, its isotropy fibres form an invariant family of ideals. Going in the other direction, if $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ is an invariant family of ideals, we define
[…]
By [8, Corollary 2.9], $I(\mathcal{K})$ is an ideal in $C^{*}(G)$. We isolate the following observation from [8].
Lemma 1.3.
Let $\mathcal{K}=(K_{x})_{x\in G^{(0)}}$ be an invariant family of ideals, and let $A\subseteq C^{*}(G)$ be a sub-$C^{*}$-algebra. Then $A\subseteq I(\mathcal{K})$ if and only if $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$.
Proof.
Assume that $A$ satisfies $A\subseteq I(\mathcal{K})$. Fix $x\in G^{(0)}$. We have that $\eta_{x}(a)\in K_{x}$ for all positive elements $a\in A^{+}$. Since $A$ is spanned by its positive elements, it follows that $\eta_{x}(a)\in K_{x}$ for all $a\in A$. Conversely, assume that $\eta_{x}(A)\subseteq K_{x}$ for all $x\in G^{(0)}$. For any $a\in A$ we have $\eta_{x}(a^{*}a)\in K_{x}$ for all $x\in G^{(0)}$, implying that $a\in I(\mathcal{K})$ as desired. ∎
2. Ideals are determined by their isotropy fibres
[…]
Remark 2.1.
[…]
Our first observation is that the assignment $I\mapsto\tilde{I}$ is injective.
Proposition 2.2.
For any ideal $I$ in $C^{*}(G)$, we have $\mathfrak{i}^{-1}(\tilde{I})=I$.
Proof.
[…]
Let $X\in\tilde{G}^{(0)}$ and write $x\coloneq\pi(X)\in G^{(0)}$. By definition, $X$ is equal to a subgroup of the isotropy group $G^{x}_{x}$. It is easy to see that the isotropy group $\tilde{G}^{X}_{X}\subseteq\tilde{G}$ is equal to the quotient $N_{X}/X$, where $N_{X}=\{g\in G^{x}_{x}:gXg^{-1}=X\}$ is the normaliser of $X$ in $G^{x}_{x}$. Denote by $E_{N_{X}}\colon C^{*}(G^{x}_{x})\to C^{*}(N_{X})$ the canonical conditional expectation and $Q_{X}\colon C^{*}(N_{X})\to C^{*}(N_{X}/X)$ the -homomorphism induced by the quotient map $N_{X}\to N_{X}/X$.
Lemma 2.3.
Let $X\in\tilde{G}^{(0)}$ and write $x\coloneq\pi(X)\in G^{(0)}$. The following diagram commutes.
$C^{*}(G)$${C^{*}(\tilde{G})}$$C^{*}(G^{x}_{x})$$C^{*}(\tilde{G}^{X}_{X})$$\scriptstyle{\mathfrak{i}}$$\scriptstyle{\eta_{x}}$$\scriptstyle{\eta_{X}}$$\scriptstyle{Q_{X}\circ E_{N_{X}}}$
Proof.
Equality is verified readily on the dense *-subalgebra $\mathscr{C}_{c}(G)$, and the lemma follows immediately. ∎
We can now describe the isotropy fibres of the ideals $\tilde{I}$ $C^{*}(\tilde{G}) in$ terms of $I$.
Proposition 2.4.
Let $I$ be an ideal in $C^{*}(G)$. Let $X\in\tilde{G}^{(0)}$ and write $x\coloneq\pi(X)\in G^{(0)}$. Then,
$\tilde{I}_{X}=\overline{Q_{X}\circ E_{N_{X}}(I_{x})}\trianglelefteq C^{*}(\tilde{G}^{X}_{X}).$
Proof.
For each $X\in\tilde{G}^{(0)}$ define the ideal $K_{X}\coloneq\overline{Q_{X}\circ E_{N_{X}}(I_{x})}\trianglelefteq C^{*}(\tilde{G}^{X}_{X})$. Let $\bm{g}\in\tilde{G}$, and observe that $\Psi_{\bm{g}}\circ Q_{s(\bm{g})}\circ E_{N_{s(\bm{g})}}=Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\circ\Psi_{g_{0}}$ for any $g_{0}\in\bm{g}$, where $\Psi_{\bm{g}}$ and $\Psi_{g_{0}}$ are as defined in subsection 1.3 (this equality can be seen using (1.1)). Since $\Psi_{\bm{g}}$ is an isomorphism, and since $(I_{x})_{x\in G^{(0)}}$ is an invariant family of ideals, it follows that
$\Psi_{\bm{g}}(K_{s(\bm{g})})$ $=\overline{\Psi_{\bm{g}}\big(Q_{s(\bm{g})}\circ E_{N_{s(\bm{g})}}(I_{s(g_{0})})\big)}=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}\big(\Psi_{g_{0}}(I_{s(g_{0})})\big)}$
$=\overline{Q_{r(\bm{g})}\circ E_{N_{r(\bm{g})}}(I_{r(g_{0})})}=K_{r(\bm{g})}.$
Hence, $\mathcal{K}\coloneq(K_{X})_{X\in\tilde{G}^{(0)}}$ is an invariant family of ideals.
By Lemma 2.3, we have $K_{X}=\overline{\eta_{X}\big(\mathfrak{i}(I)\big)}$, and therefore $K_{X}\subseteq\tilde{I}_{X}$ for all $X\in\tilde{G}^{(0)}$. By Lemma 1.3, this equality also implies that $\mathfrak{i}(I)\subseteq I(\mathcal{K})$. Hence $\tilde{I}\subseteq I(\mathcal{K})$, and another application of Lemma 1.3 gives $\tilde{I}_{X}\subseteq K_{X}$ for all $X\in\tilde{G}^{(0)}$, as desired. ∎
We now prove that ideals in $C^{*}(G)$ are determined by their isotropy fibres whenever $G$ is an amenable second-countable étale groupoid. This generalises [8, Theorem 2.10] to the setting of non-Hausdorff groupoids. Our proof utilises this theorem, which in turn uses known cases of the Effros-Hahn conjecture [17]. Therefore it is necessary for us to only consider amenable second-countable groupoids. The theorem does not hold in the non-amenable setting - counterexamples are provided in [8, Examples 2.11 & 2.12].
Theorem 2.5.
Let $G$ be an amenable second-countable (non-Hausdorff) étale groupoid and suppose $I$ and $K$ are ideals in $C^{*}(G)$. Then, $I=K$ if and only if $I_{x}=K_{x}$ for all $x\in G^{(0)}$.
Proof.
The forwards implication is trivial. Assume $I_{x}=K_{x}$ for all $x\in G^{(0)}$. Proposition 2.4 implies that $\tilde{I}_{X}=\tilde{K}_{X}$ for all $X\in\tilde{G}^{(0)}$. Since $G$ is amenable and second-countable, so is $\tilde{G}$ by [4, Theorem 6.5]. Moreover, the groupoid $\tilde{G}$ is Hausdorff, and therefore [8, Theorem 2.10] ensures that $\tilde{I}=\tilde{K}$. Finally, Proposition 2.2 gives $I=\mathfrak{i}^{-1}(\tilde{I})=\mathfrak{i}^{-1}(\tilde{K})=K$. ∎
2.1. Characterisation of density
In this subsection, we assume that $G$ is an amenable second-countable (not necessarily Hausdorff) étale groupoid, and therefore $C^{*}(G)=C^{*}_{r}(G)$ by [4, Corollary 6.10]. We denote by $J$ the singular ideal in $C^{*}_{r}(G)$. Theorem 2.5 yields the following characterisation for the density of $J\cap\mathscr{C}_{c}(G)$ in $J$.
Proposition 2.6.
Let $G$ be an amenable and second-countable étale groupoid. Then, $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ if and only if $J_{x}\cap\mathbb{C}[G^{x}_{x}]$ is dense in $J_{x}$ for all $x\in G^{(0)}$.
Proof.
Let $I=\overline{J\cap\mathscr{C}_{c}(G)}$. Since $\mathscr{C}_{c}(G)$ is dense in $C^{*}_{r}(G)$, $I$ is an ideal and $I\subseteq J$. By [16, Theorem 5.5 & Proposition 5.19], we have $\eta_{x}(J\cap\mathscr{C}_{c}(G))=J_{x}\cap\mathbb{C}[G^{x}_{x}]$. Continuity of $\eta_{x}$ then implies $J_{x}\cap\mathbb{C}[G^{x}_{x}]\subseteq\eta_{x}(I)\subseteq\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$. Since $\eta_{x}(I)=I_{x}$ is closed, it follows that $I_{x}=\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}$. Therefore, Theorem 2.5 shows that $I=J$ if and only if $\overline{J_{x}\cap\mathbb{C}[G^{x}_{x}]}=J_{x}$ for all $x\in G^{(0)}$, as desired. ∎
Remark 2.7.
The “only if” direction of Proposition 2.6 holds for all étale groupoids $G$ because $\eta_{x}(J\cap\mathscr{C}_{c}(G))\subseteq J_{x}\cap\mathbb{C}[G^{x}_{x}]$ holds trivially ($J_{x}$ here means the isotropy fibre from the reduced groupoid $C^{*}$-algebra). It can be shown using the calculation of isotropy fibres of the singular ideal in [16, Theorem 5.5] that the first example in [20] of a non-Hausdorff groupoid $G$ with $J\cap\mathscr{C}_{c}(G)$ not dense in $J$ has an isotropy fibre $J_{\epsilon}\neq\{0\}$ with $J_{\epsilon}\cap\mathbb{C}[\Gamma]=\{0\}$. Therefore, this remark may be applied to this example to provide an alternate proof of non-density.
Following [16], we now explain how the density question can be reduced to one about explicit ideals in group $C^{*}$-algebras. Let $\Gamma$ be an amenable discrete group. Let $\text{Sub}(\Gamma)$ denote the set of all subgroups in $\Gamma$. Consider the map
$\text{Sub}(\Gamma)$ $\longrightarrow\{0,1\}^{\Gamma}$
$X$ $\longmapsto\mathbbm{1}_{X}$
where $\mathbbm{1}_{X}\colon\Gamma\to\{0,1\}$ is the characteristic function, and equip $\{0,1\}^{\Gamma}$ with the product topology. We equip $\text{Sub}(\Gamma)$ with the topology induced by this map, known as the Chabauty topology, making $\text{Sub}(\Gamma)$ a Stone space (this topology agrees with the Fell topology of [14]).
For a subgroup $X\in\text{Sub}(\Gamma)$, write $\Gamma/X$ for the set of left-cosets. Let $\lambda_{\Gamma/X}\colon\Gamma\to B\big(\ell^{2}(\Gamma/X)\big)$ denote the associated quasi-regular representation as described in [3]. Concretely, we have
$\lambda_{\Gamma/X}(g)\delta_{hX}\coloneq\delta_{ghX}$ for $g\in\Gamma$ and $hX\in\Gamma/X$, where $\delta_{hX}\in\ell^{2}(\Gamma/X)$ denotes a standard basis vector. We also write $\lambda_{\Gamma/X}$ for the associated -representation on $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$.
Let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be closed in the Chabauty topology and invariant under conjugation by elements of $\Gamma$. Following [16], we let the ideal $J_{\Gamma,\mathcal{X}}$ denote the $C^{*}$-kernel of $\lambda_{\Gamma/\mathcal{X}}\coloneq\oplus_{X\in\mathcal{X}}\lambda_{\Gamma/X}$ inside $C^{*}_{r}(\Gamma)=C^{*}(\Gamma)$. Concretely, we have $J_{\Gamma,\mathcal{X}}=\bigcap_{X\in\mathcal{X}}\text{ker}(\lambda_{\Gamma/X})$.
Definition 2.8.
Let $\Gamma$ be an amenable discrete group. Define $\mathcal{D}_{\Gamma}$ to be the set of all closed conjugation-invariant sets of subgroups $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ for which $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$. We will say $\Gamma$ satisfies the Density Property if $\mathcal{D}_{\Gamma}$ contains all closed conjugation-invariant sets of subgroups.
Let $G$ be an amenable second-countable étale groupoid. For $x\in G^{(0)}$ write $\mathcal{X}(x)\coloneq\pi^{-1}(x)\cap\tilde{G}^{(0)}_{{\rm{ess}}}$, where $\pi\colon\tilde{G}^{(0)}\to G^{(0)}$ is the canonical surjection and $\tilde{G}^{(0)}_{\rm{ess}}$ is the unit space of the essential Hausdorff cover (see Subsection 1.2). The isotropy groups of $G$ are amenable, and in this case it was shown in [16, Definition 5.1 & Theorem 5.5] that $J_{x}=J_{G^{x}_{x},\mathcal{X}(x)}$ for every $x\in G^{(0)}$. The following Theorem is then immediate.
Theorem 2.9.
Let $G$ be an amenable and second-countable étale groupoid. Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ if and only if $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ for all $x\in G^{(0)}$. In particular, if all the isotropy groups $G^{x}_{x}$ satisfy the Density Property, then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Remark 2.10.
Since $J_{\Gamma,\mathcal{X}}=\{0\}$ whenever $\{e\}\in\mathcal{X}$, in the above characterisation one only needs to check whether $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ whenever $\{x\}\notin\mathcal{X}(x)$. These units $x$ are shown in [16, Corollary 5.6] to be precisely the extremely dangerous points described in [24, Proposition 1.12]. Note also that the set of units $x\in G^{(0)}$ for which $\mathcal{X}(x)\in\mathcal{D}_{G^{x}_{x}}$ is a $G$-invariant set.
It is clear that finite groups satisfy the Density Property, and hence $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for any amenable second-countable étale groupoid with finite isotropy groups.
Question 2.11.
Which discrete amenable groups satisfy the Density Property?
Assume that there exists a (discrete) countable amenable group $\Gamma$ which fails to have the Density Property, and let $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ be a closed conjugation-invariant set of subgroups for which $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is not dense in $J_{\Gamma,\mathcal{X}}$, i.e. $\mathcal{X}\notin\mathcal{D}_{\Gamma}$. Note that the trivial subgroup, $\{e\}$, is not in $\mathcal{X}$ (amenability of $\Gamma$ ensures that the quasi-regular representation $\lambda_{\Gamma/\{e\}}$ has trivial kernel). By [16, Section 6], there exists an amenable second-countable étale groupoid $G$ and a unit element $x_{0}\in G^{(0)}$ for which $G^{x_{0}}_{x_{0}}=\Gamma$ and $\mathcal{X}(x_{0})=\mathcal{X}$. It follows from Theorem 2.9 that $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$. Therefore, we have established the following.
Corollary 2.12.
The intersection $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for all amenable second-countable étale groupoids if and only if all countable amenable groups satisfy the Density Property.
3. Groups with the Density Property
In this section, we study the set $\mathcal{D}_{\Gamma}$ for amenable discrete groups $\Gamma$, and show that certain classes of groups satisfy the Density Property (see Definition 2.8). Using Theorem 2.9, we are then able to deduce that $J\cap\mathscr{C}_{c}(G)$ is dense in the singular ideal $J$ for certain classes of groupoids.
Throughout this section we work with the reduced group $C^{*}$-algebra $C^{*}_{r}(\Gamma)$ since this is canonically isomorphic to the full $C^{*}$-algebra. The set of subgroups $\text{Sub}(\Gamma)$ will always be equipped with the Chabauty topology, and given a closed and conjugation-invariant subset $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ we write $J_{\Gamma,\mathcal{X}}\coloneq\bigcap_{X\in\mathcal{X}}\text{ker}\lambda_{\Gamma/X}\trianglelefteq C^{*}_{r}(\Gamma)$ (see Subsection 2.1).
Given a subgroup $\Lambda\subseteq\Gamma$, we write $E_{\Lambda}\colon C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Lambda)$ for the canonical conditional expectation and $I_{\Lambda}\colon C^{*}_{r}(\Lambda)\to C^{*}_{r}(\Gamma)$ for the canonical inclusion. If $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ is closed and conjugation-invariant, then $\Lambda\cap\mathcal{X}\coloneq\{\Lambda\cap X:X\in\mathcal{X}\}$ is a closed and conjugation-invariant set of subgroups of $\Lambda$ by [16, Proposition 7.4]. We isolate the following observations from [16, Propositions 7.4 & 7.7].
Lemma 3.1.
Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ be a closed and conjugation-invariant set of subgroups.
  1. (i)
    For any subgroup $\Lambda\subseteq\Gamma$ we have
    $I_{\Lambda}\big(J_{\Lambda,\Lambda\cap\mathcal{X}}\big)=J_{\Gamma,\mathcal{X}}\cap I_{\Lambda}\big(C^{*}_{r}(\Lambda)\big)\quad\text{and}\quad I_{\Lambda}\big(J_{\Lambda,\Lambda\cap\mathcal{X}}\cap\mathbb{C}[\Lambda]\big)=J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Lambda]$
  2. (ii)
    Let $N$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. Then,
    $E_{N}(J_{\Gamma,\mathcal{X}})=J_{N,\mathcal{X}}\quad\text{and}\quad E_{N}(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma])=J_{N,\mathcal{X}}\cap\mathbb{C}[N].$
Proof.
For a proof of (i) see [16, Proposition 7.4]. For a proof of (ii), see the proof of [16, Proposition 7.7] (where $\Phi$ denotes $E_{N}$). ∎
We begin our study into the Density Property by showing it is preserved under countable increasing unions.
Proposition 3.2.
Suppose that an amenable discrete group $\Gamma$ is the countable increasing union of subgroups $\Lambda_{1}\subseteq\Lambda_{2}\subseteq\dots\subseteq\Gamma$. If $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ is closed, conjugation-invariant, and such that $\Lambda_{n}\cap\mathcal{X}\in\mathcal{D}_{\Lambda_{n}}$ for all $n\in\mathbb{N}$, then $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
In particular, if all $\Lambda_{n}$ have the Density Property, then so does $\Gamma$.
Proof.
For each $n\in\mathbb{N}$, write $\mathcal{X}_{n}\coloneq\Lambda_{n}\cap\mathcal{X}$. Since $\mathcal{X}_{n}\in\mathcal{D}_{\Lambda_{n}}$ for every $n\in\mathbb{N}$, Lemma 3.1 (i) implies that $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Lambda_{n}]$ is dense in $J_{\Gamma,\mathcal{X}}\cap I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ for every $n\in\mathbb{N}$. Hence, $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]=\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Lambda_{n}]$ is dense in $\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$. The countable increasing union $\bigcup_{n=1}^{\infty}I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ is dense in $C^{*}_{r}(\Gamma)$, so $\bigcup_{n}J_{\Gamma,\mathcal{X}}\cap I_{\Lambda_{n}}\big(C^{*}_{r}(\Lambda_{n})\big)$ is dense in $J_{\Gamma,\mathcal{X}}$ by [11, Lemma III.4.1], completing the proof. ∎
Next, we prove a general lemma about ideals and dense -subalgebras.
Lemma 3.3.
Let A be a $C^{*}$-algebra, $\mathcal{A}\subseteq A$ a dense -subalgebra, and $I\trianglelefteq A$ an ideal. Then $I\cap\mathcal{A}$ is dense in $I$ if and only if $I\cap\mathcal{A}$ contains a net $(u_{\beta})$ of contractions such that $\lim_{\beta}\|a-au_{\beta}\|=0$ for all $a\in I$. Moreover, $u_{\beta}$ can be chosen to be positive.
We will need the following consequence of this lemma: For any closed and conjugation-invariant set $\mathcal{X}\subseteq\text{Sub}(\Gamma)$, we have $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ contains a net of positive contractions converging strongly to the identity in $J_{\Gamma,\mathcal{X}}$.
Proof.
Assume that $I\cap\mathcal{A}$ is dense in $I$. Since $I$ is a $C^{*}$-algebra, it contains an approximate unit $(u_{\beta})$. Then $I\cap\mathcal{A}$ is a dense -subalgebra, so each $u_{\beta}$ can be approximated by a sequence of positive elements in the unit ball of $I\cap\mathcal{A}$. By a diagonalisation argument, $I\cap\mathcal{A}$ contains a net of positive elements $I$ converging strongly to the identity in $I$. Conversely, assume that $I\cap\mathcal{A}$ contains a net of contractions converging strongly to the identity in $I$. Let $a\in I$ and $\varepsilon>0$. There exist $u\in I\cap\mathcal{A}$ and $b\in\mathcal{A}$ such that $\left\|u\right\|\leq 1$, $\left\|a-au\right\|<\frac{\varepsilon}{2}$ and $\left\|a-b\right\|<\frac{\varepsilon}{2}$. Then $bu\in I\cap\mathcal{A}$ and
$\left\|a-bu\right\|\leq\left\|a-au\right\|+\left\|(a-b)u\right\|<\varepsilon.$
We now prove our first application of this lemma.
Proposition 3.4.
Let $\Gamma$ be an amenable discrete group and suppose $\mathcal{X}_{1},\dots,\mathcal{X}_{n}\subseteq\mathrm{Sub}(\Gamma)$ are closed and conjugation-invariant sets of subgroups satisfying $\mathcal{X}_{i}\in\mathcal{D}_{\Gamma}$ for all $i=1,\dots,n$. Then, $\mathcal{X}\coloneq\bigcup_{i=1}^{n}\mathcal{X}_{i}$ satisfies $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
Proof.
By definition, we have $J_{\Gamma,\mathcal{X}}=\bigcap^{n}_{i=1}J_{\Gamma,\mathcal{X}_{i}}$. For each $i=1,\dots,n$ there exists, by Lemma 3.3, a net $(u_{i,\beta})$ of positive contractions in $J_{\Gamma,\mathcal{X}_{i}}\cap\mathbb{C}[\Gamma]$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}_{i}}$ and hence in $J_{\Gamma,\mathcal{X}}\subseteq J_{\Gamma,\mathcal{X}_{i}}$. Therefore, by a diagonalisation argument, the net $(u_{\beta_{1},\beta_{2}..,\beta_{n}}):=(u_{1,\beta_{1}}\cdot u_{2,\beta_{n}}\cdot...\cdot u_{n,\beta_{n}})$ (equipped with the lexicographic partial ordering on $(\beta_{1},...,\beta_{n})$) has a subnet $(u_{\gamma})$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}}$. The net $(u_{\gamma})$ lies in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$, so $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Lemma 3.3. ∎
Next, we show the Density Property is preserved under taking quotients of groups by normal subgroups. For an amenable discrete group $\Gamma$ and a normal subgroup $N$, let $Q_{N}:C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Gamma/N)$ be the *-homomorphism induced by the quotient map $q_{N}:\Gamma\to\Gamma/N$. The map
$q^{*}_{N}:\text{Sub}(\Gamma/N)$ $\rightarrow\text{Sub}(\Gamma),\text{ }X\mapsto q_{N}^{-1}(X),$
is a continuous injection satisfying $\text{Conj}_{g}\circ q^{*}_{N}=q^{*}_{N}\circ\text{Conj}_{q_{N}(g)}$ for all $g\in\Gamma$, where $\text{Conj}_{g}$ denotes conjugation by $g$. Therefore, if $\mathcal{X}$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then $q^{*}_{N}(\mathcal{X})$ is a closed and conjugation-invariant set of subgroups of $\Gamma$.
Theorem 3.5.
Suppose $\Gamma$ is an amenable discrete group and $N$ a normal subgroup. If $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma/N)$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then
$J_{\Gamma,q_{N}^{*}(\mathcal{X})}=Q_{N}^{-1}(J_{\Gamma/N,\mathcal{X}}).$
Moreover, if $q^{*}_{N}(\mathcal{X})\in\mathcal{D}_{\Gamma}$, then $\mathcal{X}\in\mathcal{D}_{\Gamma/N}$. Consequently, if $\Gamma$ satisfies the Density Property, then so does $\Gamma/N$.
Proof.
For $X\in\mathcal{X}$, it is easy to see that $\lambda_{\Gamma/q_{N}^{-1}(X)}=\lambda_{(\Gamma/N)/\mathcal{X}}\circ Q_{N}$ and therefore $\ker(\lambda_{\Gamma/q_{N}^{-1}(X)})=Q_{N}^{-1}(\ker(\lambda_{(\Gamma/N)/\mathcal{X}}))$. Hence,
$J_{\Gamma,q^{*}_{N}(\mathcal{X})}=\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/q_{N}^{-1}(X)})=Q^{-1}_{N}\big(\bigcap_{X\in\mathcal{X}}\ker(\lambda_{(\Gamma/N)/X})\big)=Q^{-1}_{N}(J_{\Gamma/N,\mathcal{X}}).$
Now, suppose $q_{N}^{*}(\mathcal{X})\in\mathcal{D}_{\Gamma}$. By Lemma 3.3, we may choose a net of positive contractions $(u_{\lambda})\subseteq J_{\Gamma,q^{*}_{N}(\mathcal{X})}\cap\mathbb{C}[\Gamma]$ converging strongly to the identity in $J_{\Gamma,q^{*}_{N}(\mathcal{X})}$. Since $J_{\Gamma,q_{N}^{*}(\mathcal{X})}=Q_{N}^{-1}(J_{\Gamma/N,\mathcal{X}})$, $Q_{N}$ surjects $J_{\Gamma,q_{N}^{*}(\mathcal{X})}$ onto $J_{\Gamma/N,\mathcal{X}}$ and therefore $(Q_{N}(u_{\lambda}))\subseteq J_{\Gamma/N,\mathcal{X}}\cap\mathbb{C}[\Gamma/N]$ is a net of positive contractions converging strongly to the identity in $J_{\Gamma/N,\mathcal{X}}$. By Lemma 3.3, it follows that $\mathcal{X}\in\mathcal{D}_{\Gamma/N}$. ∎
We will now begin to prove the following result.
Theorem 3.6.
Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. Let $\Lambda\subseteq\Gamma$ be a subgroup such that $X\subseteq\Lambda$ for all $X\in\mathcal{X}$. Then, $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $\mathcal{X}\in\mathcal{D}_{\Lambda}$.
The following lemma is probably well-known to experts in group theory - it allows us to approximate an element $a\in C^{*}_{r}(\Gamma)$ by its restriction to finitely many cosets of a normal subgroup $N$. If $\Gamma$ is an amenable discrete group and $N$ a normal subgroup, for $gN\in\Gamma/N$, and $a\in C^{*}_{r}(\Gamma)$, define $a_{gN}\coloneq\delta_{g}\cdot E_{N}(\delta_{g^{-1}}\cdot a)\in C^{*}_{r}(\Gamma)$. Whenever $a_{gN}=0$ for all but finitely many cosets $gN$, the (finite) sum $\sum_{gN}a_{gN}=a$.
Lemma 3.7.
Let $\Gamma$ be an amenable discrete group and $N$ a normal subgroup. Choose a Følner net $(F_{i})$ in $\Gamma/N$ and define $\varphi_{i}:\Gamma/N\to\mathbb{C}$ via $gN\mapsto\frac{|gF_{i}\cap F_{i}|}{|F_{i}|}$. For any $a\in C^{*}_{r}(\Gamma)$ define
$a_{i}\coloneq\sum_{gN\in\Gamma/N}\varphi_{i}(gN)a_{gN}.$
Then, $\|a_{i}\|\leq\|a\|$ for all $i$ and the net $(a_{i})$ converges to $a$.
Proof.
[…]
The fact that $(F_{i})$ is a Følner net implies $(\tilde{\varphi}_{i}(g))$ converges to $1$ for all $g\in\Gamma$. Therefore, $(m_{i}(a))$ converges to $a$ whenever $a\in\mathbb{C}[\Gamma]$, and hence for all $a\in C^{*}_{r}(\Gamma)$ (by boundedness of $m_{i})$. ∎
Proof of Theorem 3.6.
Let $N$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. It suffices to prove the theorem in the case $\Lambda=N$, since $N$ is independent of the ambient group.
If $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$, then continuity of $E_{N}$ and Lemma 3.1 (ii) imply that $J_{N,\mathcal{X}}\cap\mathbb{C}[N]=E_{N}(J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma])$ is dense in $J_{N,\mathcal{X}}=E_{N}(J_{\Gamma,\mathcal{X}})$. This proves the “only if” direction.
Conversely, assume that $J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ is dense in $J_{N,\mathcal{X}}$. By Lemma 3.3 there exists a net of contractions $(u_{beta})\subseteq J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ converging strongly to the identity in $J_{N,\mathcal{X}}$. Since $J_{\Gamma,\mathcal{X}}\cap I_{N}\big(C^{*}_{r}(N)\big)=I_{N}\big(J_{N,\mathcal{X}}\big)$ by Lemma 3.1 (i), we can regard $(u_{\beta})$ as a net in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$.
Let $a\in J_{\Gamma,\mathcal{X}}$ and fix a Følner net $(F_{i})$ for $\Gamma/N$, with $(a_{i})$ as in Lemma 3.7. By Lemma 3.1, $a_{gN}\in J_{\Gamma,\mathcal{X}}$ for all $gN\in\Gamma/N$. Since $a_{i}$ is a finite linear combination of the $a_{gN}$, it follows that $a_{i}\in J_{\Gamma,\mathcal{X}}$ for all $i$. Moreover, since $\delta_{g^{-1}}\cdot a_{gN}\in J_{N,\mathcal{X}}$, we have $\lim_{\beta}\|a_{gN}\cdot u_{\beta}-a_{gN}\|=0$ for all $gN\in\Gamma/N$ and therefore $\lim_{\beta}\|a_{i}\cdot u_{\beta}-a_{i}\|=0$ for all $i$. Since $(a_{i})$ converges to $a$, $\|a_{i}\|\leq\|a\|$ and $\|u_{\beta}\|\leq 1$ for all $i,\beta$, it follows from the above that $\lim_{\beta}\|a\cdot u_{\beta}-a\|=0$. Lemma 3.3 then implies $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$, completing the proof. ∎
Corollary 3.8.
If $\Gamma$ is an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ is a finite collection of normal subgroups, then $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
Proof.
By Proposition 3.4, we only need to establish the case where $\mathcal{X}=\{N\}$. By Theorem 3.6, it then suffices to establish the case where $\Gamma=N$. In this case, $J_{\Gamma,\{\Gamma\}}$ is the kernel of the trivial representation $\epsilon:C^{*}_{r}(\Gamma)\to\mathbb{C}$. Let $(F_{i})$ be a Følner net for $\Gamma$ and set $v_{i}=\frac{1}{|F_{i}|}\sum_{g\in F_{i}}\delta_{g}$. For any $g\in\Gamma$, we have $\lim_{i}\|\delta_{g}\cdot v_{i}-v_{i}\|=0$, hence (by linearity and density) $\lim_{i}\|a\cdot v_{i}-\epsilon(a)v_{i}\|=0$ for all $a\in C^{*}_{r}(\Gamma)$. By Kesten’s criteria ([5, Theorem 2.6.8(8)]) we have $\|v_{i}\|=1$ for all $i$, so the above limit implies $\lim_{i}\|a\cdot v_{i}\|=|\epsilon(a)|$ for all $a\in C^{*}_{r}(\Gamma)$. Set $u_{i}=\frac{1}{2}(\delta_{e}-v_{i})$ for all $i$. Then $(u_{i})$ is a net of contractions in $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$, and $\lim_{i}\|a\cdot u_{i}-a\|=\lim_{i}\frac{1}{2}\|a\cdot v_{i}\|=\frac{1}{2}|\epsilon(a)|=0$ for any $a\in J_{\Gamma,\{\Gamma\}}$. It follows from Lemma 3.3 that $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\{\Gamma\}}$, as desired. ∎
Theorem 3.6 shows that in order to determine whether a closed and conjugation-invariant set $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ belongs to $\mathcal{D}_{\Gamma}$, it suffices to assume that $\Gamma$ is generated by the subgroups $X\in\mathcal{X}$. The next result shows that we can also assume $\mathcal{X}$ is minimal in a certain sense.
For any discrete group $\Gamma$ and $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ closed conjugation-invariant, let $\mathcal{X}_{\text{min}}$ be the collection of $X\in\mathcal{X}$ such that $X^{\prime}\subseteq X$ and $X^{\prime}\in\mathcal{X}$ implies $X=X^{\prime}$. Then, $\mathcal{X}_{\text{min}}$ is conjugation-invariant but not necessarily closed.
Theorem 3.9.
Let $\Gamma$ be a discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. Then, for every $X\in\mathcal{X}$, there is $Y\in\mathcal{X}_{\mathrm{min}}$ such that $Y\subseteq X$. Consequently, for $\Gamma$ amenable, we have $J_{\Gamma,\mathcal{X}}=J_{\Gamma,\overline{\mathcal{X}_{\mathrm{min}}}}$ and $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $\overline{\mathcal{X}_{\mathrm{min}}}\in\mathcal{D}_{\Gamma}$.
Proof.
Fix $X\in\mathcal{X}$. Let $\mathcal{Z}_{X}=\{Y\in\mathcal{X}:Y\subseteq X\}$ and (partial) order $\mathcal{Z}_{X}$ by inclusion. Let $\mathcal{C}\subseteq\mathcal{Z}_{X}$ be a chain in $\mathcal{Z}_{X}$. Then, $\mathcal{C}$ is directed, so we can view it as a net, ordered by inclusion. By compactness of $\mathcal{X}$ the net $\mathcal{C}$ has a convergent subnet $(Y_{\beta})$, and let $Y_{*}$ be its limit point. As $(Y_{\beta})$ is a subnet, for every $Y^{\prime}\in\mathcal{C}$, there is $\beta_{0}$ such that $Y_{\beta_{0}}\subseteq Y^{\prime}$. Therefore, for all $\beta\geq\beta_{0}$, we have $Y_{\beta}\subseteq Y^{\prime}$ and hence $Y_{*}\subseteq Y^{\prime}$. Hence, $Y_{*}\subseteq Y^{\prime}\subseteq X$ for all $Y^{\prime}\in\mathcal{C}$. Therefore, every chain in $\mathcal{Z}_{X}$ has a minimal element. Zorn’s lemma implies $\mathcal{Z}_{X}$ contains a minimal element $Y\in\mathcal{\mathcal{X}}$.
If $X_{1}\subseteq X_{2}$ are subgroups of $\Gamma$, then $\ker(\lambda_{\Gamma/X_{1}})\subseteq\ker(\lambda_{\Gamma/X_{2}})$ (see [2, Appendix E and F]). Therefore, $J_{\Gamma,\mathcal{X}}=\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})$. By [3, Proposition 3.3], we have
$\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\overline{\mathcal{X}_{\text{min}}}}\ker(\lambda_{\Gamma/X})$, which completes the proof. ∎
We will now show that abelian groups satisfy the Density Property, and that collections of finite subgroups always belong to $\mathcal{D}_{\Gamma}$. First, we establish a key lemma that unifies the two cases.
Lemma 3.10.
Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. If every $X\in\mathcal{X}_{\mathrm{min}}$ is finitely generated, then $\mathcal{X}_{\mathrm{min}}$ is finite.
Proof.
For each $X\in\mathcal{X}_{\text{min}}$, the set $\mathcal{Z}(X)=\{X^{\prime}\in\mathcal{X}:X\subseteq X^{\prime}\}=\bigcap^{n}_{i=1}\{X^{\prime}\in\mathcal{X}:x_{i}\in X^{\prime}\}$, where $\{x_{1},...,x_{n}\}$ is a finite generating set for $X$. Therefore, $\mathcal{Z}(X)$ is open in $\mathcal{X}$ with respect to the Chabauty topology. Since $\mathcal{Z}(X)\cap\mathcal{X}_{\text{min}}=\{X\}$ for all $X\in\mathcal{X}_{\text{min}}$, the set $\mathcal{X}_{\text{min}}$ is discrete in the compact space $\mathcal{X}$, and is therefore finite. ∎
Now, we see our result for abelian groups and collections of finite subgroups as a simple consequence of Theorems 3.6, Theorem 3.9, Corollary 3.8 and Lemma 3.10.
Theorem 3.11.
Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. If either
  1. (I)
    $\Gamma$ is countable and abelian, or
  2. (II)
    every $X\in\mathcal{X}$ is finite,
then $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
Proof.
Let $\Gamma$ be a countable abelian group. By Proposition 3.2, it suffices to assume that $\Gamma$ is finitely generated. Then, every subgroup of $\Gamma$ is also finitely generated, so Lemma 3.10 implies that $\mathcal{X}_{\text{min}}$ is finite. Since $\Gamma$ is abelian, each $X\in\mathcal{X}_{\text{min}}$ is normal, and hence Corollary 3.8 implies $\mathcal{X}_{\text{min}}\in\mathcal{D}_{\Gamma}$. Theorem 3.9 then implies $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
In the case where every $X\in\mathcal{X}$ is finite, we have that $\mathcal{X}_{\text{min}}$ is finite (by Lemma 3.10) and each $X\in\mathcal{X}_{\text{min}}$ is finite. It follows from [16, Lemma 7.9] (see also the proof of [4, Theorem 4.7]) that the normal subgroup $N$ generated by all $X\in\mathcal{X}_{\text{min}}$ is finite. Since $C^{*}_{r}(N)=\mathbb{C}[N]$, it holds trivially that $\mathcal{X}_{\text{min}}\in\mathcal{D}_{N}$, so Theorems 3.6 and 3.9 imply $\mathcal{X}\in\mathcal{D}_{\Gamma}.$
The following Corollary is immediate by Theorems 2.9 and 3.11.
Corollary 3.12.
Let $G$ be an amenable and second-countable étale groupoid, and let $\tilde{G}$ denote the Hausdorff cover groupoid. Assume for every $x\in G^{(0)}$ one of the following holds.
  1. (I)
    The isotropy group $G_{x}^{x}$ is abelian.
  2. (II)
    The subgroups $X\in\mathcal{X}(x)$ are finite.
Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Note that condition (II) holds whenever all elements of $\tilde{G}^{(0)}_{\rm{ess}}$ are finite. The class of groupoids described in Corollary 3.12 (II) contains all amenable and second-countable étale groupoids with finite isotropy groups, and all groupoids of contracting self-similar groups (see [4, Corollary 7.13]). More generally, condition (II) is satisfied whenever the unit space of $G$ has finite source and range fibres. Therefore, within the class of amenable and second-countable étale groupoids, the class described in Corollary 3.12 (II) is larger than the class described in [4, Theorem 4.7].
4. Ample Groupoids
The main result of this section is that the algebraic singular ideal $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ for any ample groupoid. This reduces the question of whether $J_{\mathbb{C}}$ is dense in $J$ to the corresponding question for $J\cap\mathscr{C}_{c}(G)$. In order to prove this result we will need two preliminary lemmas.
Lemma 4.1.
Let $G$ be an étale groupoid, and let $f\in\mathscr{C}_{c}(G)$. Let $x\in G^{(0)}$, and assume that $f(g)=0$ for all $g\in G_{x}$. Then $f$ is continuous at each $g\in G_{x}$.
Proof.
Let $g\in G_{x}$ and let $(g_{\beta})$ be a net in $G$ converging to $g$. In order to prove that $\lim_{\beta}f(g_{\beta})=0$, it suffices to find a subnet $(g_{\gamma})$ satisfying $\lim_{\gamma}f(g_{\gamma})=0$ (since $(g_{\beta})$ is arbitrary). There exists a subnet $(g_{\gamma})$ for which the net $\iota(g_{\gamma})$ converges in $\tilde{G}$ (i.e. in the Fell topology). Let $\bm{g}\in\tilde{G}$ denote the limit. By (1.2) we have $\mathfrak{i}(f)(\bm{g})=\sum_{g\in\bm{g}}f(g)=0$ since $\bm{g}\subseteq G_{x}$. Moreover, $f(g_{\gamma})=\mathfrak{i}(f)\big(\iota(g_{\gamma})\big)\to\mathfrak{i}(f)(\bm{g})$ by continuity of $\mathfrak{i}(f)$. This completes the proof. ∎
Observations similar to the following lemma have appeared in both [4, Theorem 4.7] and [16, Proposition 5.19]. Our proof uses the same techniques as those developed in the aforementioned papers.
Lemma 4.2.
Let $G$ be an étale groupoid. Fix $f\in J\cap\mathscr{C}_{c}(G)$ and $x_{0}\in G^{(0)}$. Let $g_{1},\dotsc,g_{n}\in G_{x_{0}}$ be all points in $G_{x_{0}}$ on which $f$ is non-zero, and let $U_{1},\dotsc,U_{n}$ be open bisections such that $U_{i}$ contains $g_{i}$ for each $1\leq i\leq n$. Set $V\coloneq\bigcap_{i=1}^{n}s(U_{i})$. Then there exists an open subset $W\subseteq V$ containing $x_{0}$ such that
$f^{\psi}\coloneq\sum_{i=1}^{n}f(g_{i})(\psi\circ s|_{U_{i}})$ (4.1)
satisfies $f^{\psi}\in J$ for any $\psi\in C_{c}(W)$.
Proof.
Without loss of generality, we assume $s(U_{i})=V$ for all $1\leq i\leq n$. For each open set $W\subseteq V$, let $S^{W}_{i}\coloneq U_{i}\cap s^{-1}(W)$ for $1\leq i\leq n$. For $I\subseteq\{1,\dotsc,n\}$, set $\check{S}^{W}_{I}\coloneq(\bigcap_{i\in I}S^{W}_{i})\setminus(\bigcup_{j\notin I}S^{W}_{j})$ as well as $W_{I}\coloneq\text{int}\Big(\overline{s(\check{S}^{W}_{I})}^{G^{(0)}}\Big)$, where $\overline{(\,\cdot\,)}^{G^{(0)}}$ denotes closure in $G^{(0)}$. Our first claim is that there exists an open set $W\subseteq V$ containing $x_{0}$ such that
$W_{I}\neq\emptyset\text{ implies that }x_{0}\in\overline{W_{I}}^{G^{(0)}}\text{ for all }I\subseteq\{1,\dotsc,n\}.$ (4.2)
If $W=V$ satisfies (4.2), then we are done. Otherwise, choose a regular open subset $W$ of $G^{(0)}$ (i.e., $W=\text{int}\big(\overline{W}^{G^{(0)}}\big)$) such that $x_{0}\in W\subseteq V$ and $W\cap\overline{V_{I}}^{G^{(0)}}=\emptyset$ for all $I\subseteq\{1,\dotsc,n\}$ with $V_{I}\neq\emptyset$ and $x_{0}\notin\overline{V_{I}}^{G^{(0)}}$. Note that $\check{S}^{W}_{I}=\check{S}^{V}_{I}\cap s^{-1}(W)$, and therefore $s(\check{S}^{W}_{I})=s(\check{S}^{V}_{I})\cap W$. Since $W$ is open, we have $W_{I}=V_{I}\cap W$, and therefore $\overline{W_{I}}^{G^{(0)}}\cap W=\overline{V_{I}}^{G^{(0)}}\cap W$. Also $W_{I}=\text{int}\Big(\overline{s(\check{S}^{W}_{I})}^{G^{(0)}}\Big)\subseteq\text{int}\Big(\overline{s(\check{S}^{V}_{I})}^{G^{(0)}}\cap\overline{W}^{G^{(0)}}\Big)\subseteq V_{I}\cap\text{int}\big(\overline{W}^{G^{(0)}}\big)=V_{I}\cap W=\emptyset$ whenever $x_{0}\notin\overline{V_{I}}^{G^{(0)}}$. This completes the proof of the claim.
Define $\mathcal{I}\coloneq\{I\subseteq\{1,\dotsc,n\}\colon W_{I}\neq\emptyset\}$. Our second claim is that
$\sum_{i\in I}f(g_{i})=0\text{ for all }I\in\mathcal{I}.$ (4.3)
By [4, Lemma 4.1(ii)], $\overline{s({\rm supp}^{\circ}(f))}^{G^{(0)}}$ has empty interior. Let $I\in\mathcal{I}$, and set $\check{W}_{I}\coloneq W_{I}\setminus\overline{s({\rm supp}^{\circ}(f))}^{G^{(0)}}$. Then, $\check{W}_{I}$ is open and dense in $W_{I}$. Therefore $x_{0}$ lies in the closure of $\check{W}_{I}$ in $G^{(0)}$ by (4.2). As $s(\check{S}_{I})$ is dense in $W_{I}$ and because $\check{W}_{I}$ is open, we conclude that there is a net $(x_{\beta})$ in $\check{W}_{I}\cap s(\check{S}_{I})$ converging to $x_{0}$ in $G^{(0)}$. Let $g_{\beta}$ be the unique element of $\check{S}_{I}$ such that $s(g_{\beta})=x_{\beta}$. By passing to a subnet if necessary, we may assume that $(g_{\beta})$ converges to some $\bm{g}$ in $\tilde{G}$ (i.e., in the Fell topology). By construction, we have that $\bm{g}\cap\{g_{1},\dotsc,g_{n}\}=\{g_{i}\colon i\in I\}$. Therefore,
$\lim_{\beta}f(g_{\beta})=\lim_{\beta}\mathfrak{i}(f)(\iota(g_{\beta}))=\mathfrak{i}(f)(\bm{g})=\sum_{i\in I}f(g_{i}),$
where the third equality used (1.2). Now $s(g_{\beta})\in\check{W}_{I}$ implies that $g_{\beta}\notin{\rm supp}^{\circ}(f)$, so that $f(g_{\beta})=0$. We conclude that $\sum_{i\in I}f(g_{i})=0$, proving (4.3).
Take arbitrary $\psi\in C_{c}(W)$, and let $f^{\psi}$ be as in (4.1). We show that $f^{\psi}\in J$. For $I\in\mathcal{I}$ and $g\in\check{S}^{W}_{I}$ we have $f^{\psi}(g)=\sum_{i\in I}f(g_{i})\psi(s(g))=0$ using (4.3). It follows that ${\rm supp}^{\circ}(f^{\psi})\subseteq\bigcup_{I\,\in\,\mathcal{P}(\{1,\dotsc,n\})\setminus\mathcal{I}}\check{S}^{W}_{I}$ and therefore $s({\rm supp}^{\circ}(f^{\psi}))\subseteq\bigcup_{I\,\in\,\mathcal{P}(\{1,\dotsc,n\})\setminus\mathcal{I}}s(\check{S}^{W}_{I})$, where $\mathcal{P}$ denotes power set. Hence $s({\rm supp}^{\circ}(f^{\psi}))$ has empty interior because it is contained in a finite union of nowhere dense sets. This shows that $f^{\psi}\in J$, as desired. ∎
Remark 4.3.
Let $f\in J$, $x_{0}\in G^{(0)}$ and assume that there are only finitely many points $g_{1},\dotsc,g_{n}$ in the isotropy group $G^{x_{0}}_{x_{0}}$ on which $f$ non-zero. Let $U_{1},\dotsc,U_{n},V$ be as in the statement of Lemma 4.2. If in addition we assume that the isotropy group $G^{x_{0}}_{x_{0}}$ is amenable, then the conclusion of Lemma 4.2 still holds. That is, there exists an open set $W\subseteq V$ containing $x_{0}$ such that $f^{\psi}\in J$ for any $\psi\in C_{c}(W)$, where $f^{\psi}$ is defined as in (4.1). If the isotropy group $G^{x_{0}}_{x_{0}}$ is non-amenable, then there may exist no such $W$.
We now prove the main result of this section. Recall that an ample groupoid is an étale groupoid with a basis of compact open bisections. If $G$ is an ample groupoid, the Steinberg algebra $\mathbb{C}G$ over the complex numbers is defined in [26] as
$\mathbb{C}G\coloneq\text{span}\left\{\mathbbm{1}_{U}\colon U\subseteq G\text{ a compact open bisection}\right\}$
The algebraic singular ideal is the intersection $J_{\mathbb{C}}\coloneq J\cap\mathbb{C}G$. Since $\mathbb{C}G\subseteq\mathscr{C}_{c}(G)$ we have $J_{\mathbb{C}}\subseteq J\cap\mathscr{C}_{c}(G)$. For $f\in\mathscr{C}_{c}(G)$, define
$\left\|f\right\|_{I}\coloneq\max\left\{\sup_{x\in G^{(0)}}\sum_{g\in G_{x}}\left|f(g)\right|,\sup_{x\in G^{(0)}}\sum_{g\in G^{x}}\left|f(g)\right|\right\}.$
Theorem 4.4.
Let $G$ be an ample groupoid. Then $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to the norm $\left\|\cdot\right\|_{I}$.
The norm $\left\|\cdot\right\|_{I}$ dominates the full norm $\left\|\cdot\right\|_{C^{*}(G)}$ and hence any $C^{*}$-norm. Therefore, $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
Proof.
Let $f\in J\cap\mathscr{C}_{c}(G)$. Take a compact set $K\subseteq G$ with ${\rm supp}^{\circ}(f)$ contained in its interior. Whenever $\dot{f}\in\mathscr{C}_{c}(G)$ satisfies ${\rm supp}^{\circ}(\dot{f})\subseteq K$ we have $\left\|\dot{f}\right\|_{I}\leq C_{K}\left\|\dot{f}\right\|_{\infty}$ where $\left\|\cdot\right\|_{\infty}$ denotes the supremum norm, and $C_{K}$ is a constant depending only on $K$. Therefore, it suffices to find, for each $\varepsilon>0$, a function $f^{\prime}\in J_{\mathbb{C}}$ satisfying ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$.
Fix $\varepsilon>0$. Take $x\in s(K)$, and let $g_{1},\dotsc,g_{m}\in G_{x}$ be all points in $G_{x}$ on which $f$ is non-zero. By Lemma 4.2, there exists a compact open neighbourhood $U_{x}\subseteq G^{(0)}$ of $x$, and compact open bisections $U_{1},\dotsc,U_{m}\subseteq K$ satisfying the following: $g_{i}\in U_{i}$ and $s(U_{i})=U_{x}$ for all $i=1,\dotsc,m$, and the function $f^{x}$ defined by
$f^{x}\coloneq\sum_{i=1}^{m}f(g_{i})\mathbbm{1}_{U_{i}}$
lies in $J_{\mathbb{C}}$. Clearly ${\rm supp}^{\circ}(f^{x})\subseteq K$. We have $f|_{G_{x}}=f^{x}|_{G_{x}}$ and ${\rm supp}^{\circ}(f-f^{x})\subseteq K$, a compact set. Therefore, by Lemma 4.1, there exists a compact open neighbourhood $W_{x}\subseteq U_{x}$ of $x$ such that $\left|f(g)-f^{x}(g)\right|\leq\varepsilon$ whenever $s(g)\in W_{x}$. The sets $\{W_{x}\}_{x\in s(K)}$ form an open cover for the compact set $s(K)$, so select a finite subcover $W_{x_{1}},\dotsc,W_{x_{n}}$. Removing intersections if necessary, we may assume that the $W_{x_{j}}$ are pairwise disjoint (this is possible since the $W_{x_{i}}$ are contained in the Hausdorff space $G^{(0)}$, and are thus clopen in $G^{(0)}$). Define
$f^{\prime}\coloneq\sum_{j=1}^{n}(\mathbbm{1}_{W_{x_{j}}}\circ s)f^{x_{j}}.$
Then $f^{\prime}\in J_{\mathbb{C}}$ satisfies ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$, as required. ∎
We apply the previous theorem to the class of groupoids arising from contracting self-similar groups.
Corollary 4.5.
Let $G$ be the groupoid arising from the action of a contracting self-similar group on a finite alphabet. Then $J_{\mathbb{C}}$ is dense in $J$.
Proof.
By Theorem 4.4 it suffices to check that $J\cap\mathscr{C}_{c}(G)$ is dense in $J$. By [23] and [15] the groupoid $G$ is second-countable and amenable. Moreover, [4, Corollary 7.13] implies that every element of $\tilde{G}^{(0)}$ is finite. The result follows by Corollary 3.12 (II). ∎